OpenStudy (anonymous):

What is the probability of rolling a sum of 8 on at least one of two rolls of a pair of number cubes? possible answers: 5/8 335/1296 25/1296 4/81 not sure how to solve this!?

jimthompson5910 (jim_thompson5910):

P(At least one sum of 8) = 1 - P(No sums of 8) P(At least one sum of 8) = 1 - (1 - 5/36)^2 P(At least one sum of 8) = 1 - (31/36)^2 P(At least one sum of 8) = 1 - 961/1296 P(At least one sum of 8) = 335/1296 So the answer is choice B

OpenStudy (anonymous):

Thank youuuuu!!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

I love you